OpenStudy (anonymous):

How to evaluate z^-1/3 * z^3/4?

OpenStudy (solomonzelman):

use the same technique. \(\Large\color{blue}{ \bf z^{-1/3} \times z^{3/4}=z ^{-1/3~~+~~3/4} }\)

OpenStudy (solomonzelman):

multiply the top and bottom of first fraction times 4, and in the second fraction times 3 to add them...

OpenStudy (anonymous):

The different denominator stumped me. Okay!

OpenStudy (solomonzelman):

\[-1/3=-4/12~~~~~~~~~~~~~~3/4=9/12\]

OpenStudy (solomonzelman):

\[So,~~~-1/3+3/4~~=~~-4/12+9/12=(-4+9)/12=5/12\]